# Lorenzo Ciardo

### v1

@@ 1 Introduction
The algebraic characterisation of the power of local consistency [14, 6] was one of the fundamental milestones of the algebraic theory of constraint satisfaction problems (CSPs) [36, 15], and a key step towards the positive resolution of Feder–Vardi’s Dichotomy Conjecture [29] obtained, independently, by Bulatov [16] and Zhuk [54]. Fix a relational structure $\mathbf{A}$ (consisting of a finite domain $A$ and a finite list of relations over $A$). The constraint satisfaction problem parameterised by $\mathbf{A}$ (in short, $\operatorname{CSP}(\mathbf{A})$) is the following computational problem: Given an instance structure $\mathbf{X}$ that admits a homomorphism (i.e., a relation-preserving map) to $\mathbf{A}$, find an explicit such homomorphism. [^1] We denote the presence of a homomorphism from $\mathbf{X}$ to $\mathbf{A}$ by the expression $\mathbf{X}\to\mathbf{A}$. Suitably choosing the template $\mathbf{A}$ allows for casting as a CSP many diverse problems, including graph colouring (if $\mathbf{A}$ is a clique $\mathbf{K}_{c}$), graph homomorphism (if $\mathbf{A}$ is an undirected graph), linear equations over a finite field (if $\mathbf{A}$ encodes the affine spaces over the field), and several variants of the Boolean satisfiability problem such as Horn Sat, $k$-in-$n$ Sat, and Not-All-Equal Sat.
@@
Part of the algebraic machinery developed in the decades of research on CSPs can be transferred to PCSPs. In particular, polymorphisms have a natural promise version (homomorphisms $\mathbf{A}^{n}\to\mathbf{B}$) and, just like for CSPs, they completely determine the complexity of PCSPs, as established in [12, 5]. Yet, the CSP tools appear to lose a substantial part of their sharpness when they are adapted to the promise world. The ultimate reason for this weakening is that the common, universal-algebraic core underpinning many distinct CSP phenomena is less structured for PCSPs. In particular, PCSP polymorphisms do not compose, which makes it impossible to obtain new polymorphic identities from old ones. A clear example of this phenomenon is the break of the “WNU – local consistency” bond: While the local-consistency algorithm can be applied to solve PCSPs and, thus, the notion of width is well defined in the promise setting, the presence of WNU polymorphisms of all arities does not guarantee bounded width for PCSPs, as established in [1]. For example, the PCSP parameterised by the structures encoding $2$-in-$4$ Sat and Not-All-Equal Sat has WNU polymorphisms of all arities although its width is unbounded — and, in fact, linear; see later. [^4] Unless P$=$NP, WNU polymorphisms do not even guarantee polynomial-time tractability (via any algorithm): For instance, $\operatorname{PCSP}(\mathbf{K}_{c},\mathbf{K}_{2c})$ has WNUs of all arities; yet, it was shown in [43] that the problem is NP-hard for $c\geq 6$. In particular, this makes it infeasible to investigate the width of promise problems such as approximate graph homomorphism using the polymorphic tools currently available.
@@
We adopt a different, structural approach to study the power of local consistency. To illustrate it, we start with a simple example — $\operatorname{CSP}(\mathbf{C}_{n})$, where $\mathbf{C}_{n}$ is the undirected cycle of length $n$. In this case, the parity of $n$ determines the width of the problem: It is well known and not hard to check that $\operatorname{CSP}(\mathbf{C}_{n})$ has width $3$ if $n$ is even, while it has unbounded width if $n$ is odd. Can we link these two different width behaviours to a separating structural property depending on the parity of the cycle length?
@@
We have seen that a (P)CSP has bounded width if it is solved by a constant level of the local-consistency algorithm — in which case, the running time of the algorithm is polynomial in the size of the instance. (P)CSPs of linear width lie at the opposite end of the width spectrum: Not even $\epsilon n$ levels of the local-consistency algorithm are enough to solve these problems (where $n$ is the size of the instance and $\epsilon$ is a positive constant). [^7] In particular, every PCSP of linear width has unbounded width. For non-promise CSPs, the converse also holds: Every CSP of unbounded width has linear width [2]. For PCSPs, the analogous statement is not known to hold. Classic examples of CSPs of linear width are systems of linear equations [29, 2] (see also [30, 52, 20] for other lower bounds on systems of linear equations). Our main result identifies aperiodicity as a sufficient condition for having linear — and, thus, unbounded — width. [^8] Put differently, Theorem 1 gives a necessary condition for not having linear width, and thus in particular a necessary condition for having bounded width.
@@
Let $\mathbf{A}\to\mathbf{B}$ be relational structures such that $\mathbf{A}$ is aperiodic and $\mathbf{B}$ is loopless. [^9] A structure $\mathbf{B}$ contains a loop if there exists an element $b\in B$ such that every relation of $\mathbf{B}$ contains the constant tuple $(b,\dots,b)$ of suitable arity. Otherwise, we say that $\mathbf{B}$ is loopless. Then, $\operatorname{PCSP}(\mathbf{A},\mathbf{B})$ has linear width. [^10] Our result holds both for the search version of $\operatorname{PCSP}(\mathbf{A},\mathbf{B})$ defined above and for its decision version — distinguish whether an instance $\mathbf{X}$ is homomorphic to $\mathbf{A}$ or is not even homomorphic to $\mathbf{B}$.
Observe that $n$ levels of the local-consistency algorithm always suffice for instances of size $n$. Hence, the linear-level lower bound in Theorem 1 is tight (up to optimising the constant $\epsilon$). An immediate consequence of the result above is the complete characterisation of the power of local consistency for the approximate graph homomorphism problem.
@@
Let $\mathbf{A}\to\mathbf{B}$ be undirected graphs, and suppose that $\mathbf{A}$ is non-bipartite and $\mathbf{B}$ is loopless. Then, $\operatorname{PCSP}(\mathbf{A},\mathbf{B})$ has linear width.111111 It is well known that $\operatorname{PCSP}(\mathbf{A},\mathbf{B})$ has bounded width if $\mathbf{A}$ is a bipartite graph or $\mathbf{B}$ has a loop [29]. Note that, in the statement of Corollary 2, we are implicitly identifying an undirected graph with a symmetric digraph containing both orientations of each edge.
@@
Our work is inspired by the recent paper [1] by Atserias–Dalmau, which gives a sufficient condition for a PCSP to have linear width. [^12] While the result in [1] is stated in terms of sublinear, as opposed to linear, width, the proof yields a linear-level, optimal lower bound. In the language of the current paper, that condition requires that any pair of vertices should be connected by oriented walks of length $2$; thus, it corresponds to the case $\tau=2$ of our aperiodicity notion. Atserias–Dalmau’s condition is satisfied, in particular, by cliques, thus resulting in a linear-level lower bound against local consistency for the solution of approximate graph colouring. Extending the lower bound to approximate graph homomorphism requires dealing with graphs having arbitrarily high mixing time and, thus, it needs the full power of our Theorem 1. Some of the machinery of [1] is used in this paper as a black box (in particular, a result on the sparsity of random hypergraphs obtained via Chernoff bounds). However, the transition from $\tau=2$ to arbitrarily high $\tau$ requires rather different proof ideas. In particular, the concept of $\tau$-fibrosity that we introduce in this work to study the number of long, isolated chains of hyperedges in sparse hypergraphs turns out to behave differently for $\tau\leq 3$ and for $\tau\geq 4$: In the latter case, large fibrosity is not implied by sparsity alone, and bounds on the length of the shortest cycles in hypergraphs need to be established via a probabilistic analysis.
A recent line of work established lower bounds against relaxations of approximate graph colouring [22, 23] and homomorphism [25] based on linear and semidefinite programming as well as linear-equation solvers. NP-hardness of the former problem was proved in [31] conditional to Khot’s $d$-to-$1$ Conjecture [39]. Brakensiek–Guruswami conjectured that approximate graph homomorphism should be tractable in polynomial time if $\mathbf{A}$ is bipartite or $\mathbf{B}$ has a loop, and NP-hard otherwise [12]. Therefore, our Corollary 2 is consistent with — and supports — the conjectured complexity dichotomy for this problem.
@@ 2.1 Aperiodicity
A relational structure $\mathbf{A}$ is aperiodic if ${\mathbf{A}^{\operatorname{mon}}}$ is aperiodic; in this case, the mixing time of $\mathbf{A}$ is defined to be the mixing time of ${\mathbf{A}^{\operatorname{mon}}}$. [^16] As mentioned in the Introduction, the condition for linear width in [1, Thm. 3] is equivalent to the structure being aperiodic with mixing time $2$.
If $\mathbf{A}$ is a digraph, each row of a $\tau$-pattern $\lambda$ is either the pair $(1,2)$ (the “forward” direction) or the pair $(2,1)$ (the “backward” direction). Thus, in this case, Definition 3 asks for the existence of some $\tau$ such that every two vertices $a,b$ are connected by a walk of length $\tau$, oriented in every possible way. It turns out that considering two orientations only (“directed” and “alternating”) is enough — thus recovering the definition encountered in the Introduction for the digraph case. In fact, via matrix-theoretic results (in particular, a classical theorem by Wielandt on the maximum index of primitivity of primitive matrices), we can make this condition even weaker. This also allows upper bounding the mixing time of an aperiodic digraph.
@@ 2.2 Fibrosity
For an integer $r\geq 2$, an $r$-uniform hypergraph $\mathcal{H}=(V(\mathcal{H}),E(\mathcal{H}))$ consists of a nonempty set $V(\mathcal{H})$ of vertices and a set $E(\mathcal{H})$ of hyperedges, where a hyperedge is a set of vertices of cardinality exactly $r$. Henceforth, we shall let $r$ be fixed, and we shall refer to $r$-uniform hypergraphs simply as hypergraphs. We denote the number of vertices and the number of hyperedges of $\mathcal{H}$ by $n_{\mathcal{H}}$ and $m_{\mathcal{H}}$, respectively. Two hyperedges are adjacent if they have a nonempty intersection; a vertex is incident to a hyperedge if it belongs to it. For $v\in V(\mathcal{H})$, $\deg_{\mathcal{H}}(v)$ is the degree of $v$; i.e., the number of hyperedges of $\mathcal{H}$ to which $v$ is incident. We say that a hyperedge $e$ is a link if exactly two vertices in $e$ have degree $2$, and all of the other $r-2$ vertices have degree $1$. A set $f$ of distinct links is called a fiber if the elements of $f$ can be labelled as $e_{1},\dots,e_{\tau}$, for some $\tau\in\mathbb{N}$, in a way that $e_{i}$ is adjacent to $e_{i+1}$ for each $i\in[\tau-1]$. If we need to stress that $f$ has cardinality $\tau$, we refer to it as a $\tau$-fiber. The $\tau$-fibrosity of $\mathcal{H}$, which we denote by $\varphi^{\tau}_{\mathcal{H}}$, is the maximum cardinality of a set of mutually disjoint $\tau$-fibers in $\mathcal{H}$.
@@
For $\tau\in\mathbb{N}$ and $1<\beta\in\mathbb{R}$, let $\mathcal{H}$ be a hereditarily $\beta$-sparse hypergraph of girth at least $\tau$ having no isolated vertices. [^18] An isolated vertex is a vertex of degree 00. Then, $\varphi^{\tau}_{\mathcal{H}}+\pi_{\mathcal{H}}>\left(\frac{1}{10r\tau}-\beta+1\right)n_{\mathcal{H}}$.
@@ 2.5 Epilogue
A natural question, at this point, is whether the sufficient condition for linear width in Corollary 15 is also necessary. The answer turns out to be negative, even in the CSP setting. Indeed, it is known that there exists an oriented tree T whose corresponding CSP has unbounded width [11]. By [2] (see Footnote 7), $\operatorname{CSP}(\textbf{T})$ has in fact linear width. Suppose that there exists some aperiodic digraph $\mathbf{A}^{\prime}$ such that $\mathbf{A}^{\prime}\to\textbf{T}$. It is not hard to check (see Proposition 25) that this implies the existence of some aperiodic induced substructure $\textbf{T}^{\prime}$ of T. This is a contradiction, as $\textbf{T}^{\prime}$ is an oriented forest and, thus, no directed walk in $\textbf{T}^{\prime}$ connects a vertex to itself. Another natural question is whether there exist PCSPs having intermediate width; i.e., super-constant but nonlinear. In the two cases of non-promise CSPs and binary symmetric PCSPs no such problems exist, as follows from [2] and from Corollary 2, respectively. Does the width dichotomy extend to the entire class of PCSPs?
@@ 3 The local-consistency algorithm
In this preliminary section, we formally describe the local-consistency algorithm, following [41, 8]. Let $\mathbf{X}$ and $\mathbf{A}$ be two relational structures having the same signature $\sigma$. We say that a structure $\mathbf{Y}$ of signature $\sigma$ is a substructure of $\mathbf{X}$ if $Y\subseteq X$ and $R^{\mathbf{Y}}\subseteq R^{\mathbf{X}}$ for each $R\in\sigma$. If, in particular, $R^{\mathbf{Y}}=R^{\mathbf{X}}\cap Y^{\operatorname{ar}(R)}$ for each $R\in\sigma$, we say that $\mathbf{Y}$ is an induced substructure of $\mathbf{X}$. (We define subhypergraphs and induced subhypergraphs analogously.) A partial homomorphism from $\mathbf{X}$ to $\mathbf{A}$ is either a homomorphism from some induced substructure of $\mathbf{X}$ to $\mathbf{A}$, or the empty mapping from $\emptyset$ to $A$. Take an integer $\kappa\in\mathbb{N}$, and consider a nonempty family $\mathscr{F}$ of partial homomorphisms from $\mathbf{X}$ to $\mathbf{A}$. Given a function $f$, let $\operatorname{dom}(f)$ denote its domain. We say that $\mathscr{F}$ is a $\kappa$-strategy for $\mathbf{X}$ and $\mathbf{A}$ if it is closed under restrictions (i.e., for every $f\in\mathscr{F}$ and $Y\subseteq\operatorname{dom}(f)$, the restriction of $f$ to $Y$ is in $\mathscr{F}$) and has the extension property up to $\kappa$ (i.e., for every $f\in\mathscr{F}$ with $|\operatorname{dom}(f)|<\kappa$ and every $x\in X\setminus\operatorname{dom}(f)$, there exists some $f^{\prime}\in\mathscr{F}$ that extends $f$ and has domain $\operatorname{dom}(f)\cup\{x\}$). Testing for the existence of a $\kappa$-strategy for $\mathbf{X}$ and $\mathbf{A}$ can be performed in time polynomial in $(n_{\mathbf{X}}+n_{\mathbf{A}})^{\kappa}$ through the so-called local-consistency algorithm, which starts with all partial homomorphisms from $\mathbf{X}$ to $\mathbf{A}$ with domain size at most $\kappa$, and iteratively discards those that do not satisfy the two conditions above, until a fixed point is reached. If the resulting family of partial homomorphisms is nonempty, it must be a $\kappa$-strategy; in this case, we write $\operatorname{LC}^{\kappa}(\mathbf{X},\mathbf{A})=\textsc{Yes}$. Otherwise, we are guaranteed that no $\kappa$-strategy exists, and we write $\operatorname{LC}^{\kappa}(\mathbf{X},\mathbf{A})=\textsc{No}$. [^21] It shall be convenient to let $\operatorname{LC}^{0}(\mathbf{X},\mathbf{A})=\textsc{Yes}$ for each $\mathbf{X}$ and $\mathbf{A}$.

### v2

@@ 1 Introduction
The algebraic characterisation of the power of local consistency [14, 6] was one of the fundamental milestones of the algebraic theory of constraint satisfaction problems (CSPs) [40, 15], and a key step towards the positive resolution of Feder–Vardi’s Dichotomy Conjecture [30] obtained, independently, by Bulatov [16] and Zhuk [58]. Fix a relational structure $\mathbf{A}$ (consisting of a finite domain $A$ and a finite list of relations over $A$). The constraint satisfaction problem parameterised by $\mathbf{A}$ (in short, $\operatorname{CSP}(\mathbf{A})$) is the following computational problem: Given an instance structure $\mathbf{X}$ that admits a homomorphism (i.e., a relation-preserving map) to $\mathbf{A}$, find an explicit such homomorphism. We denote the presence of a homomorphism from $\mathbf{X}$ to $\mathbf{A}$ by the expression $\mathbf{X}\to\mathbf{A}$. Suitably choosing the template $\mathbf{A}$ allows for casting as a CSP many diverse problems, including graph colouring (if $\mathbf{A}$ is a clique $\mathbf{K}_{c}$), graph homomorphism (if $\mathbf{A}$ is an undirected graph), linear equations over a finite field (if $\mathbf{A}$ encodes the affine spaces over the field), and several variants of the Boolean satisfiability problem such as Horn Sat, $k$-in-$n$ Sat, and Not-All-Equal Sat.
@@
Part of the algebraic machinery developed in the decades of research on CSPs can be transferred to PCSPs. In particular, polymorphisms have a natural promise version (homomorphisms $\mathbf{A}^{n}\to\mathbf{B}$) and, just like for CSPs, they completely determine the complexity of PCSPs, as established in [12, 5]. Yet, the CSP tools appear to lose a substantial part of their sharpness when they are adapted to the promise world. The ultimate reason for this weakening is that the common, universal-algebraic core underpinning many distinct CSP phenomena is less structured for PCSPs. In particular, PCSP polymorphisms do not compose, which makes it impossible to obtain new polymorphic identities from old ones. A clear example of this phenomenon is the break of the “WNU – local consistency” bond: While the local-consistency algorithm can be applied to solve PCSPs and, thus, the notion of width is well defined in the promise setting, the presence of WNU polymorphisms of all arities does not guarantee bounded width for PCSPs, as established in [1]. For example, the PCSP parameterised by the structures encoding $2$-in-$4$ Sat and Not-All-Equal Sat has WNU polymorphisms of all arities although its width is unbounded—and, in fact, linear; see later. Unless P$=$NP, WNU polymorphisms do not even guarantee polynomial-time tractability (via any algorithm): For instance, $\operatorname{PCSP}(\mathbf{K}_{c},\mathbf{K}_{2c})$ has WNUs of all arities; yet, it was shown in [47] that the problem is NP-hard for $c\geq 6$. In particular, this makes it infeasible to investigate the width of promise problems such as approximate graph homomorphism using the polymorphic tools currently available.
@@
We adopt a different, structural approach to study the power of local consistency. To illustrate it, we start with a simple example—CSP($\operatorname{CSP}(\mathbf{C}_{n})$, where $\mathbf{C}_{n}$ is the undirected cycle of length $n$. In this case, the parity of $n$ determines the width of the problem: It is well known and not hard to check that $\operatorname{CSP}(\mathbf{C}_{n})$ has width $3$ if $n$ is even, while it has unbounded width if $n$ is odd. Can we link these two different width behaviours to a separating structural property depending on the parity of the cycle length?
@@
We have seen that a (P)CSP has bounded width if it is solved by a constant level of the local-consistency algorithm—in which case, the running time of the algorithm is polynomial in the size of the instance. (P)CSPs of linear width lie at the opposite end of the width spectrum: Not even $\epsilon n$ levels of the local-consistency algorithm are enough to solve these problems (where $n$ is the size of the instance and $\epsilon$ is a positive constant). [^5] In particular, every PCSP of linear width has unbounded width. For non-promise CSPs, the converse also holds: Every CSP of unbounded width has linear width [2]. For PCSPs, the analogous statement is not known to hold. Classic examples of CSPs of linear width are systems of linear equations [30, 2] (see also [34, 56, 20] for other lower bounds on systems of linear equations). Our main result identifies aperiodicity as a sufficient condition for having linear—and, thus, unbounded—width.
@@
Let $\mathbf{A}\to\mathbf{B}$ be relational structures such that $\mathbf{A}$ is aperiodic and $\mathbf{B}$ is loopless. [^6] A structure $\mathbf{B}$ contains a loop if there exists an element $b\in B$ such that every relation of $\mathbf{B}$ contains the constant tuple $(b,\dots,b)$ of suitable arity. Otherwise, we say that $\mathbf{B}$ is loopless. Then, $\operatorname{PCSP}(\mathbf{A},\mathbf{B})$ has linear width.
Our result holds both for the search version of $\operatorname{PCSP}(\mathbf{A},\mathbf{B})$ defined above and for its decision version—distinguish whether an instance $\mathbf{X}$ is homomorphic to $\mathbf{A}$ or is not even homomorphic to $\mathbf{B}$. Observe that $n$ levels of the local-consistency algorithm always suffice for instances of size $n$. Hence, the linear-level lower bound in Theorem 1 is tight (up to optimising the constant $\epsilon$). An immediate consequence of the result above is the complete characterisation of the power of local consistency for the approximate graph homomorphism problem.
@@
Let $\mathbf{A}\to\mathbf{B}$ be undirected graphs, and suppose that $\mathbf{A}$ is non-bipartite and $\mathbf{B}$ is loopless. Then, $\operatorname{PCSP}(\mathbf{A},\mathbf{B})$ has linear width.777 It is well known that $\operatorname{PCSP}(\mathbf{A},\mathbf{B})$ has bounded width if $\mathbf{A}$ is a bipartite graph or $\mathbf{B}$ has a loop [30]. Note that, in the statement of Corollary 2, we are implicitly identifying an undirected graph with a symmetric digraph containing both orientations of each edge.
@@
Our work is inspired by the recent paper [1] by Atserias–Dalmau, which gives a sufficient condition for a PCSP to have linear width. [^8] While the result in [1] is stated in terms of sublinear, as opposed to linear, width, the proof yields a linear-level, optimal lower bound. In the language of the current paper, that condition requires that any pair of vertices should be connected by oriented walks of length $2$; thus, it corresponds to the case $\tau=2$ of our aperiodicity notion. Atserias–Dalmau’s condition is satisfied, in particular, by cliques, thus resulting in a linear-level lower bound against local consistency for the solution of approximate graph colouring. Extending the lower bound to approximate graph homomorphism requires dealing with graphs having arbitrarily high mixing time and, thus, it needs the full power of our Theorem 1. Some of the machinery of [1] is used in this paper as a black box (in particular, a result on the sparsity of random hypergraphs obtained via Chernoff bounds). However, switching from $\tau=2$ to arbitrarily high $\tau$ requires rather different proof ideas. In particular, the concept of $\tau$-fibrosity that we introduce in this work to study the number of long, isolated chains of hyperedges in sparse hypergraphs turns out to behave differently for $\tau\leq 3$ and for $\tau\geq 4$: In the latter case, large fibrosity is not implied by sparsity alone, and bounds on the length of the shortest cycles in hypergraphs need to be established via a probabilistic analysis.
The occurrence of long paths and cycles in sparse random hypergraphs and the corresponding phase transitions have recently been investigated in various different settings via probabilistic-combinatorial methods, see [26, 33, 32, 31]. A recent line of work established lower bounds against relaxations of approximate graph colouring [22, 23] and homomorphism [25] based on linear and semidefinite programming as well as linear-equation solvers. NP-hardness of the former problem was proved in [35] conditional to Khot’s $d$-to-$1$ Conjecture [43]. Brakensiek–Guruswami conjectured that approximate graph homomorphism should be tractable in polynomial time if $\mathbf{A}$ is bipartite or $\mathbf{B}$ has a loop, and NP-hard otherwise [12]. Therefore, our Corollary 2 is consistent with—and supports—the conjectured complexity dichotomy for this problem.
@@ 2.1 Aperiodicity
A relational structure $\mathbf{A}$ is aperiodic if ${\mathbf{A}^{\operatorname{mon}}}$ is aperiodic; in this case, the mixing time of $\mathbf{A}$ is defined to be the mixing time of ${\mathbf{A}^{\operatorname{mon}}}$.
As mentioned in the Introduction, the condition for linear width in [1, Thm. 3] is equivalent to the structure being aperiodic with mixing time $2$. If $\mathbf{A}$ is a digraph, each row of a $\tau$-pattern $\lambda$ is either the pair $(1,2)$ (the “forward” direction) or the pair $(2,1)$ (the “backward” direction). Thus, in this case, Definition 3 asks for the existence of some $\tau$ such that every two vertices $a,b$ are connected by a walk of length $\tau$, oriented in every possible way. It turns out that considering two orientations only (“directed” and “alternating”) is enough—thus recovering the definition encountered in the Introduction for the digraph case. In fact, via matrix-theoretic results (in particular, a classical theorem by Wielandt on the maximum index of primitivity of primitive matrices), we can make this condition even weaker. This also allows upper bounding the mixing time of an aperiodic digraph.
@@ 2.2 Fibrosity
For an integer $r\geq 2$, an $r$-uniform hypergraph $\mathcal{H}=(V(\mathcal{H}),E(\mathcal{H}))$ consists of a nonempty set $V(\mathcal{H})$ of vertices and a set $E(\mathcal{H})$ of hyperedges, where a hyperedge is a set of vertices of cardinality exactly $r$. Henceforth, we shall let $r$ be fixed, and we shall refer to $r$-uniform hypergraphs simply as hypergraphs. We denote the number of vertices and the number of hyperedges of $\mathcal{H}$ by $n_{\mathcal{H}}$ and $m_{\mathcal{H}}$, respectively. Two hyperedges are adjacent if they have a nonempty intersection; a vertex is incident to a hyperedge if it belongs to it. For $v\in V(\mathcal{H})$, $\deg_{\mathcal{H}}(v)$ is the degree of $v$; i.e., the number of hyperedges of $\mathcal{H}$ to which $v$ is incident. A vertex is isolated if its degree is 00. We say that a hyperedge $e$ is a link if exactly two vertices in $e$ have degree $2$, and all of the other $r-2$ vertices have degree $1$. A set $f$ of distinct links is called a fiber if the elements of $f$ can be labelled as $e_{1},\dots,e_{\tau}$, for some $\tau\in\mathbb{N}$, in a way that $e_{i}$ is adjacent to $e_{i+1}$ for each $i\in[\tau-1]$. If we need to stress that $f$ has cardinality $\tau$, we refer to it as a $\tau$-fiber. The $\tau$-fibrosity of $\mathcal{H}$, which we denote by $\varphi^{\tau}_{\mathcal{H}}$, is the maximum cardinality of a set of mutually disjoint $\tau$-fibers in $\mathcal{H}$.
@@
For $\tau\in\mathbb{N}$ and $1<\beta\in\mathbb{R}$, let $\mathcal{H}$ be a hereditarily $\beta$-sparse hypergraph of girth at least $\tau$ having no isolated vertices. Then, $\varphi^{\tau}_{\mathcal{H}}+\pi_{\mathcal{H}}>\left(\frac{1}{10r\tau}-\beta+1\right)n_{\mathcal{H}}$.
@@ 2.5 Epilogue
A natural question, at this point, is whether the sufficient condition for linear width in Corollary 15 is also necessary. The answer turns out to be negative, even in the CSP setting. Indeed, it is known that there exists an oriented tree T whose corresponding CSP has unbounded width [11]. By [2] (see Footnote 5), $\operatorname{CSP}(\textbf{T})$ has in fact linear width. Suppose that there exists some aperiodic digraph $\mathbf{A}^{\prime}$ such that $\mathbf{A}^{\prime}\to\textbf{T}$. It is not hard to check (see Proposition 25) that this implies the existence of some aperiodic induced substructure $\textbf{T}^{\prime}$ of T. This is a contradiction, as $\textbf{T}^{\prime}$ is an oriented forest and, thus, no directed walk in $\textbf{T}^{\prime}$ connects a vertex to itself. Another natural question is whether there exist PCSPs having intermediate width; i.e., super-constant but nonlinear. In the two cases of non-promise CSPs and binary symmetric PCSPs no such problems exist, as follows from [2] and from Corollary 2, respectively. Does the width dichotomy extend to the entire class of PCSPs?
@@ 3 The local-consistency algorithm
In this preliminary section, we formally describe the local-consistency algorithm, following [45, 8]. Let $\mathbf{X}$ and $\mathbf{A}$ be two relational structures having the same signature $\sigma$. We say that a structure $\mathbf{Y}$ of signature $\sigma$ is a substructure of $\mathbf{X}$ if $Y\subseteq X$ and $R^{\mathbf{Y}}\subseteq R^{\mathbf{X}}$ for each $R\in\sigma$. If, in particular, $R^{\mathbf{Y}}=R^{\mathbf{X}}\cap Y^{\operatorname{ar}(R)}$ for each $R\in\sigma$, we say that $\mathbf{Y}$ is an induced substructure of $\mathbf{X}$. (We define subhypergraphs and induced subhypergraphs analogously.) A partial homomorphism from $\mathbf{X}$ to $\mathbf{A}$ is either a homomorphism from some induced substructure of $\mathbf{X}$ to $\mathbf{A}$, or the empty mapping from $\emptyset$ to $A$. Take an integer $\kappa\in\mathbb{N}$, and consider a nonempty family $\mathscr{F}$ of partial homomorphisms from $\mathbf{X}$ to $\mathbf{A}$. Given a function $f$, let $\operatorname{dom}(f)$ denote its domain. We say that $\mathscr{F}$ is a $\kappa$-strategy for $\mathbf{X}$ and $\mathbf{A}$ if it is closed under restrictions (i.e., for every $f\in\mathscr{F}$ and $Y\subseteq\operatorname{dom}(f)$, the restriction of $f$ to $Y$ is in $\mathscr{F}$) and has the extension property up to $\kappa$ (i.e., for every $f\in\mathscr{F}$ with $|\operatorname{dom}(f)|<\kappa$ and every $x\in X\setminus\operatorname{dom}(f)$, there exists some $f^{\prime}\in\mathscr{F}$ that extends $f$ and has domain $\operatorname{dom}(f)\cup\{x\}$). Testing for the existence of a $\kappa$-strategy for $\mathbf{X}$ and $\mathbf{A}$ can be performed in time polynomial in $(n_{\mathbf{X}}+n_{\mathbf{A}})^{\kappa}$ through the so-called local-consistency algorithm, which starts with all partial homomorphisms from $\mathbf{X}$ to $\mathbf{A}$ with domain size at most $\kappa$, and iteratively discards those that do not satisfy the two conditions above, until a fixed point is reached. If the resulting family of partial homomorphisms is nonempty, it must be a $\kappa$-strategy; in this case, we write $\operatorname{LC}^{\kappa}(\mathbf{X},\mathbf{A})=\textsc{Yes}$. Otherwise, we are guaranteed that no $\kappa$-strategy exists, and we write $\operatorname{LC}^{\kappa}(\mathbf{X},\mathbf{A})=\textsc{No}$. It shall be convenient to let $\operatorname{LC}^{0}(\mathbf{X},\mathbf{A})=\textsc{Yes}$ for each $\mathbf{X}$ and $\mathbf{A}$.